My father is 24 years older than me, and 12 years ago he was double my age. How old am i now?
Answer by stanbon(75887)

My father is 24 years older than me, and 12 years ago he was double my age. How old am i now?
------
Equations:
f = m + 24
f-12 = 2(m-12)
----
Substitute for "f" and solve for "m":
m+24-12 = 2m-24
m+12 = 2m-24
m = 36 (You are 36 years old)
===============
Cheers,
Stan H.